OpenStudy (lilcathyx):

I'll give a medal.. Point X is located at (2, 4), and Y is located at (−6, 10). Find the y value of the point that lies halfway between X and Y.

OpenStudy (lilcathyx):

A. 2 B. 7 C. -3 D. -8

OpenStudy (lilcathyx):

\[m = x1 + x2 \over 2 \] and \[m= y1 + y2 \over 2\]

OpenStudy (lilcathyx):

2 + (-6)/2 = -4/2 = -2 y= 4 + 10/2 = 40/2 = 20

OpenStudy (lilcathyx):

but none of those are the answer choice :/ im stuck

OpenStudy (lilcathyx):

@Hero @paki @ganeshie8 help ?

OpenStudy (anonymous):

D. -8

OpenStudy (lilcathyx):

May i ask why is it -8 ? just so if i get another question like this i know how to resolve it.. @-StudyShark-